# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v5

@@ 1.7.2. The symplectic case
Remarkably, the situation in the case of the coideal subalgebras quantizing $U(\mathfrak{sp}_{2n})$ and $U(\widehat{\mathfrak{sp}}_{2n})$ is quite different. Classification results were obtained in [Mol06, GM10], using the RTT presentation. In particular, in the affine case, there exists a classification of irreducibles in terms of Drinfeld polynomials. The nicer behaviour of the symplectic case is, to a large extent, due to the fact that these coideal subalgebras contain some of the Cartan elements $K_{i}$. This facilitates the formulation of a highest weight theory compatible with restriction.
Presently, a Drinfeld-type presentation is only available for aKM QSP of split type. This corresponds to Satake diagrams with no arrows or black dots, and in type $\mathsf{A}$ leads to coideal subalgebras quantizing $U(\widehat{\mathfrak{so}}_{n})$. However, it is natural to expect that a Drinfeld presentation should exist for all aKM QSP, even though the technical complexity of the proofs is likely to increase with the complexity of the Satake diagrams. We also expect that, once the Drinfeld presentation becomes available in new cases, Theorem 1.1 and methods used to establish it will quite straightforwardly generalize to these new cases. In particular, in the symplectic case, corresponding to the Satake diagram with an alternating colouring, the analogue of our theorem should have more direct representation theoretic consequences, allowing us to recover the classification from [GM10] from the point of view of the Drinfeld presentation.
@@ 1.7.5. Factorization of the universal K-matrix
Originally, the notion of a $q$-character for quantum affine algebras was formulated in terms of the universal R-matrix [FR99]. Its connection to spectra of the operators $\phi_{i,k}^{\pm}$ was established using the Levendorsky–Soibelman factorization formula [KT94, LSS93, Dam98]. In the QSP case, we are proceeding in the opposite direction, i.e., we propose to define $q$-characters in terms of the spectra of the operators $\Theta_{i,k}$. We expect that an equivalent formulation in terms of the universal K-matrix from [AV22b] is also possible. In particular, there should exist an analogue of the Levendorsky–Soibelman formula for the universal K-matrix. In the finite case, similar factorization results were already established in [WZ23].
@@ 2. The $q$-Onsager algebra
Throughout this paper, we assume that $q\in{\mathbb{C}}^{\times}$ is not a root of unity.
@@ 2.1. Quantum affine $\mathfrak{sl}_{2}$
(2.1) $\Delta(e_{i}^{+})=e_{i}^{+}\otimes 1+K_{i}\otimes e_{i},\quad\Delta(e_{i}^{-})=e_{i}^{-}\otimes K_{i}^{-1}+1\otimes e_{i}^{-},\quad\Delta(K_{i}^{\pm 1})=K_{i}^{\pm 1}\otimes K_{i}^{\pm 1}.$
@@ 2.2. The quantum loop algebra
(2.10) $\boldsymbol{\psi}(z)=\sum_{k=0}^{\infty}\psi_{k}z^{k}=$ $\ K\exp\left((q-q^{-1})\sum_{k=1}^{\infty}h_{k}z^{k}\right),$
(2.11) $\boldsymbol{\phi}(z)=\sum_{k=0}^{\infty}\phi_{-k}z^{-k}=$ $\ K^{-1}\exp\left(-(q-q^{-1})\sum_{k=1}^{\infty}h_{-k}z^{-k}\right).$
@@
(2.12) $\psi_{0}=K,\quad\phi_{0}=K^{-1},\quad\psi_{1}=(q-q^{-1})Kh_{1},\quad\phi_{-1}=-(q-q^{-1})K^{-1}h_{-1}.$
@@ 2.3. The $q$-Onsager algebra
Let $\mathcal{O}_{q}^{\mathbf{c}}$ be the corresponding central reduction, i.e., the quotient of $\mathcal{O}_{q}$ by the ideal generated by $\mathfrak{c}_{i}-c_{i}$. We make no distinction in the notation between elements of $\mathcal{O}_{q}$ and their images in $\mathcal{O}_{q}^{\mathbf{c}}$.
@@ 2.6. One-dimensional modules
Let $\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}=\eta_{\mathbf{c},\mathbf{s}}(\mathcal{O}_{q})$. By [DK19, Sec. 3.5], the character $\xi_{\mathbf{c},\mathbf{s}}$ induces an algebra isomorphism
@@
such that $\eta_{\mathbf{c},\mathbf{s}}=\phi_{\mathbf{c},\mathbf{s}}\circ\eta_{\mathbf{c},0}$. Explicitly, $\phi_{\mathbf{c},\mathbf{s}}$ is given by $\phi_{\mathbf{c},\mathbf{s}}=(\widetilde{\xi}_{\mathbf{c},\mathbf{s}}\otimes\operatorname{id})\circ\Delta|_{\mathcal{O}_{q}^{\mathbf{c},0}}$, where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}=\xi_{\mathbf{c},\mathbf{s}}\circ(\eta_{\mathbf{c},0}|_{\mathcal{O}_{q}^{\mathbf{c},0}})^{-1}$.
@@ 2.7. $C$-Duality
    (2.25) $M=\bigoplus_{\bm{\lambda}=(\lambda_{s})_{s\in{\mathbb{Z}}_{\geq 0}}}M_{\bm{\lambda}},\qquad M_{\bf\lambda}=\{v\in M\mid\exists p\ \forall s\in{\mathbb{Z}}_{\geq 0}:(\Theta_{s}-\lambda_{s})^{p}\cdot x=0\},$
@@
If $M_{\bm{\lambda}}\neq\{0\}$, we call $\bm{\lambda}$ a weight of $M$. In the context of quantum affine algebras, this is usually referred to as an $\ell$-weight or a pseudo-weight (see, e.g., [Her05, HJ12]).
@@ 3. Rationality
$\overline{H}_{1}=[2]^{-1}H_{1},\quad\overline{h}_{\pm 1}=[2]^{-1}h_{\pm 1}.$
@@
Let $V$ be a module in $\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q})$. Set $C=q^{4}c_{0}c_{1}$. Then, for every weight $\bm{\lambda}$ of $V$:
@@
    The generating series $\boldsymbol{A}_{\pm}{(z)}\in{\operatorname{End}(V_{\bm{\lambda}})}[\negthinspace[z^{\pm 1}]\negthinspace]$ are the expansions of the same rational function [^1] In the later sections we will typically use the same notation for rational functions and their series expansions where no confusion can arise.
@@ 4. Factorization of $\boldsymbol{\Theta}(z)$
The goal of this section is to prove our first main result, namely a factorization theorem for $\boldsymbol{\Theta}(z)$ (Theorem 4.9), and interpret it in terms of Drinfeld rational functions (Corollary 5.5). The factorization theorem establishes a connection between the Drinfeld presentation of the $q$-Onsager algebra and the Drinfeld presentation of $U_{q}L\mathfrak{sl}_{2}$. It is key to understanding restricted representations.
More precisely, we are able to express the image of the series $\boldsymbol{\Theta}(z)$ in $U_{q}L\mathfrak{sl}_{2}$ as a product of shifted series $\boldsymbol{\phi}(z)$ and $\boldsymbol{\psi}(z)$, modulo $UX_{+}$. The proof relies on an explicit computation of the homogeneous components of low degree of the images of $\mathcal{O}_{q}$-generators. The main step in the argument is the calculation of $\boldsymbol{A}^{(1)}_{+}(z)$ in §4.3.
@@ 4.1. Auxiliary relations
(4.2) $(z-q^{-2}w)\boldsymbol{\phi}(z)\boldsymbol{x}^{+}(w)=(q^{-2}z-w)\boldsymbol{x}^{+}(w)\boldsymbol{\phi}(z).$
@@
$x^{+}_{k}\boldsymbol{\phi}(z^{-1})=q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
$[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k}.$
@@
(4.8) $x^{+}_{k}\boldsymbol{\phi}(z^{-1})=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)\sum_{r\geq 0}\sum_{s=1}^{r}q^{2(s-1)}\phi_{s-r}x^{+}_{k-s}z^{r}$
(4.9) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\sum_{r\geq 0}\sum_{s=1}^{r}\phi_{s-r}q^{2(s-k)}x^{+}_{k-s}z^{r-k}$
(4.10) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq k-1}(q^{-2}z^{-1})$
(4.11) $=$ $\ q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
(4.12) $[x^{+}_{k},\boldsymbol{\phi}(z^{-1})]=$ $\ (q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}$
(4.13) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(x^{+}_{k}+x^{+}_{k-1}z)$
(4.14) $=$ $\ \boldsymbol{\phi}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k},$
@@
$(z-q^{-2}w)\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{-}(w^{-1})=(q^{-2}z-w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}(z^{-1}).$
@@ 4.2. Basic implications of the Lu–Wang relations
Next, we compute $A_{r}^{(-1)}$. Relation (2.17) (with $m=1$) implies that
@@ 4.3. $\boldsymbol{A}^{(1)}_{+}(z)$ calculation
(4.30) $\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
@@
(4.36) $x^{+}_{0}x^{+}_{-1}x^{-}_{-r}=$ $\ x^{+}_{0}(x^{-}_{-r}x^{+}_{-1}-(q-q^{-1})^{-1}\phi_{-r-1})$
(4.37) $=$ $\ -(q-q^{-1})^{-1}((\phi_{-r}-\delta_{r=0}\psi_{0})x^{+}_{-1}+x^{+}_{0}\phi_{-r-1})+x^{-}_{-r}x^{+}_{0}x^{+}_{-1},$
(4.38) $x^{+}_{-1}x^{+}_{0}x^{-}_{-r}=$ $\ x^{+}_{-1}(x^{-}_{-r}x^{+}_{0}-(q-q^{-1})^{-1}(\phi_{-r}-\delta_{r=0}\psi_{0}))$
(4.39) $=$ $\ -(q-q^{-1})^{-1}(\phi_{-r-1}x^{+}_{0}+x^{+}_{-1}(\phi_{-r}-\delta_{r=0}\psi_{0}))+x^{-}_{-r}x^{+}_{-1}x^{+}_{0}$
(4.40) $=$ $\ -q^{-2}(q-q^{-1})^{-1}((\phi_{-r}-\delta_{r=0}\psi_{0})x^{+}_{-1}+x^{+}_{0}\phi_{-r-1})+x^{-}_{-r}x^{+}_{-1}x^{+}_{0},$
@@
(4.43) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ c_{1}Kx^{+}_{-1}-q^{-2}c_{1}\sum_{r\geq 0}(q^{4}\phi_{-r-1}x^{+}_{0}+q^{6}\phi_{-r}x^{+}_{-1})z^{r}$
(4.44) $\ -c_{1}(q^{2}-q^{-2})\sum_{r\geq 0}\sum_{s=2}^{r+1}q^{2s}\phi_{s-r-1}x^{+}_{-s}z^{r}.$
(4.45) $=$ $\ c_{1}Kx^{+}_{-1}+c_{1}q^{2}K^{-1}x^{+}_{0}z^{-1}-c_{1}q^{2}\boldsymbol{\phi}(z^{-1})x^{+}_{0}z^{-1}$
(4.46) $\ -c_{1}q^{4}\boldsymbol{\phi}(z^{-1})x^{+}_{-1}-c_{1}(q^{2}-q^{-2})\boldsymbol{\phi}(z^{-1})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})z^{-1}.$
@@
(4.47) $\Omega\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}(z^{-1})\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
Dividing both sides by $\Omega$ and using the fact that $\Omega$ commutes with $\boldsymbol{\phi}(z^{-1})$ completes the proof. ∎
@@
so that $\boldsymbol{A}^{(1)}_{+}(z)=\boldsymbol{\phi}(z^{-1})\mathbb{A}$.

### v6

@@ 1.7.2. The symplectic case
Remarkably, the situation in the case of the QSP coideal subalgebras of type $\mathsf{AII}$ (also known as symplectic twisted $q$-Yangians) is quite different. Classification results were obtained in [Mol06, GM10], using the RTT presentation. In particular, there exists a classification of irreducibles in terms of Drinfeld polynomials. The nicer behaviour of the symplectic case is, to a large extent, due to the fact that these coideal subalgebras contain some of the Cartan elements $K_{i}$. This facilitates the formulation of a highest weight theory compatible with restriction.
Presently, a Drinfeld-type presentation is only available for aKM QSP of quasi-split type [BK20, LW21, Zha22, LWZ23, LWZ24], corresponding to Satake diagrams with no black dots. However, it is natural to expect that a Drinfeld presentation should exist for all aKM QSP, even though the technical complexity of the proofs is likely to increase with the complexity of the Satake diagrams. We also expect that, once the Drinfeld presentation becomes available in new cases, Theorem 1.1 and methods used to establish it will quite straightforwardly generalize to these new cases. In particular, in the symplectic case, corresponding to the Satake diagram with an alternating colouring, the analogue of our theorem should have more direct representation theoretic consequences, allowing us to recover the classification from [GM10] from the point of view of the Drinfeld presentation.
@@ 1.7.5. Factorization of the universal K-matrix
Originally, the notion of a $q$-character for quantum affine algebras was formulated in terms of the universal R-matrix [FR99]. Its connection to spectra of the operators $\phi_{i,k}^{\pm}$ was established using an infinite product factorization formula [KT94, LSS93, Dam98]. In the QSP case, we are proceeding in the opposite direction, i.e., we propose to define $q$-characters in terms of the spectra of the operators $\Theta_{i,k}$. We expect that an equivalent formulation in terms of the universal K-matrix from [AV22b] is also possible. In particular, there should exist an analogue of the Levendorsky–Soibelman formula for the universal K-matrix. In the finite case, similar factorization results were already established in [DK19, WZ23].
@@ 2. The $q$-Onsager algebra
Throughout this paper, we assume that $q\in{\mathbb{C}}^{\times}$ is not a root of unity. We use the standard notation for quantum numbers, i.e.,
[…]
for $k,l\in{\mathbb{Z}}_{+}$, and $[-k]=-[k]$. We also set $[a,b]_{q}=ab-qba$, as usual.
@@ 2.1. Quantum affine $\mathfrak{sl}_{2}$
(2.1) $\Delta(e_{i}^{+})=e_{i}^{+}\otimes 1+K_{i}\otimes e_{i}^{+},\quad\Delta(e_{i}^{-})=e_{i}^{-}\otimes K_{i}^{-1}+1\otimes e_{i}^{-},\quad\Delta(K_{i}^{\pm 1})=K_{i}^{\pm 1}\otimes K_{i}^{\pm 1}.$
@@ 2.2. The quantum loop algebra
(2.10) $\boldsymbol{\phi}^{+}(z)=\sum_{k=0}^{\infty}\phi^{+}_{k}z^{k}=$ $\ K\exp\left((q-q^{-1})\sum_{k=1}^{\infty}h_{k}z^{k}\right),$
(2.11) $\boldsymbol{\phi}^{-}(z)=\sum_{k=0}^{\infty}\phi^{-}_{-k}z^{-k}=$ $\ K^{-1}\exp\left(-(q-q^{-1})\sum_{k=1}^{\infty}h_{-k}z^{-k}\right).$
@@
(2.12) $\phi^{+}_{0}=K,\quad\phi^{-}_{0}=K^{-1},\quad\phi^{+}_{1}=(q-q^{-1})Kh_{1},\quad\phi^{-}_{-1}=-(q-q^{-1})K^{-1}h_{-1}.$
@@ 2.3. The $q$-Onsager algebra
Let $\mathcal{O}_{q}^{\mathbf{c}}$ be the corresponding central reduction, i.e., the quotient of $\mathcal{O}_{q}$ by the ideal generated by $\mathfrak{c}_{i}-c_{i}$. We make no distinction in the notation between elements of $\mathcal{O}_{q}$ and their images in $\mathcal{O}_{q}^{\mathbf{c}}$. It is known that algebras $\mathcal{O}_{q}^{\mathbf{c}}$ and $\mathcal{O}_{q}^{\mathbf{c}^{\prime}}$ are isomorphic for any choices of parameters $\mathbf{c},\mathbf{c}^{\prime}$ (see, e.g., [Kol14, Proposition 9.2] or [Wat21, Lemma 2.5.1]).
@@ 2.6. One-dimensional modules
Let $\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}=\eta_{\mathbf{c},\mathbf{s}}(\mathcal{O}_{q}^{\mathbf{c}})$. By [DK19, Sec. 3.5], the character $\xi_{\mathbf{c},\mathbf{s}}$ induces an algebra isomorphism
@@
such that $\eta_{\mathbf{c},\mathbf{s}}=\phi_{\mathbf{c},\mathbf{s}}\circ\eta_{\mathbf{c},0}$. Explicitly, $\phi_{\mathbf{c},\mathbf{s}}$ is given by $\phi_{\mathbf{c},\mathbf{s}}=(\widetilde{\xi}_{\mathbf{c},\mathbf{s}}\otimes\operatorname{id})\circ\Delta|_{\mathcal{O}_{q}^{\mathbf{c},0}}$, where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}=\xi_{\mathbf{c},\mathbf{s}}\circ(\eta^{\prime}_{\mathbf{c},0})^{-1}$, and $\eta^{\prime}_{\mathbf{c},0}$ denotes $\eta_{\mathbf{c},0}$ with codomain restricted to its image $\mathcal{O}_{q}^{\mathbf{c},0}$.
@@ 2.7. $C$-Duality
    (2.26) $M=\bigoplus_{\bm{\lambda}=(\lambda_{s})_{s\in{\mathbb{Z}}_{\geq 0}}}M_{\bm{\lambda}},\qquad M_{\bm{\lambda}}=\{v\in M\mid\exists p\ \forall s\in{\mathbb{Z}}_{\geq 0}:(\Theta_{s}-\lambda_{s})^{p}\cdot v=0\},$
@@
  4. (d)
    For each $\bm{\lambda}$, the vector space generated by $A_{r}\cdot M_{\bm{\lambda}}$ $(r\in{\mathbb{Z}})$ is finite-dimensional.
If $M_{\bm{\lambda}}\neq\{0\}$, we call $\bm{\lambda}$ a weight of $M$. In the context of quantum affine algebras, this is usually referred to as an $\ell$-weight or a pseudo-weight (see, e.g., [Her05, HJ12]). If $\bm{\lambda}$ and $\bm{\mu}$ are weights, we write $\bm{\lambda}\to\bm{\mu}$ if there exists $v\in M_{\bm{\lambda}}$ and $r\in{\mathbb{Z}}$ such that $A_{r}\cdot v$ has a non-trivial projection onto $M_{\bm{\mu}}$.
@@ 3. Rationality
$\overline{H}_{1}=[2]^{-1}H_{1},\qquad\overline{h}_{\pm 1}=[2]^{-1}h_{\pm 1},\qquad\operatorname{ad}_{\overline{H}_{1}}(-)=[\overline{H}_{1},-].$
@@
Let $V$ be a module in $\operatorname{Mod^{fdw}_{\mathbf{c}}}(\mathcal{O}_{q})$. Set
$C=q^{4}c_{0}c_{1}.$
Then, for every weight $\bm{\lambda}$ of $V$:
@@
    The generating series $\boldsymbol{A}_{\pm}{(z)}\in{\operatorname{Hom}(V_{\bm{\lambda}},\bigoplus_{\bm{\lambda}\to\bm{\mu}}V_{\bm{\mu}})}[\negthinspace[z^{\pm 1}]\negthinspace]$ are the expansions of the same rational function [^2] In the later sections we will typically use the same notation for rational functions and their series expansions where no confusion can arise.
@@ 4. Factorization of $\boldsymbol{\Theta}(z)$
The goal of this section is to prove our first main result, namely a factorization theorem for $\boldsymbol{\Theta}(z)$ (Theorem 4.9), and interpret it in terms of Drinfeld rational functions (Theorem 5.5). The factorization theorem establishes a connection between the Drinfeld presentation of the $q$-Onsager algebra and the Drinfeld presentation of $U_{q}L\mathfrak{sl}_{2}$. It is key to understanding restricted representations.
More precisely, we are able to express the image of the series $\boldsymbol{\Theta}(z)$ in $U_{q}L\mathfrak{sl}_{2}$ as a product of shifted series $\boldsymbol{\phi}^{+}(z)$ and $\boldsymbol{\phi}^{-}(z)$, modulo $U_{+}$. The proof relies on an explicit computation of the homogeneous components of low degree of the images of $\mathcal{O}_{q}$-generators. The main step in the argument is the calculation of $\boldsymbol{A}^{(1)}_{+}(z)$ in §4.3.
@@ 4.1. Auxiliary relations
(4.2) $(z-q^{-2}w)\boldsymbol{\phi}^{-}(z)\boldsymbol{x}^{+}(w)=(q^{-2}z-w)\boldsymbol{x}^{+}(w)\boldsymbol{\phi}^{-}(z).$
@@
$x^{+}_{k}\boldsymbol{\phi}^{-}(z^{-1})=q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
$[x^{+}_{k},\boldsymbol{\phi}^{-}(z^{-1})]=\boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k}.$
@@
(4.8) $x^{+}_{k}\boldsymbol{\phi}^{-}(z^{-1})=$ $\ q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{4}-1)\sum_{r\geq 0}\sum_{s=1}^{r}q^{2(s-1)}\phi^{-}_{s-r}x^{+}_{k-s}z^{r}$
(4.9) $=$ $\ q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\sum_{r\geq 0}\sum_{s=1}^{r}\phi^{-}_{s-r}q^{2(s-k)}x^{+}_{k-s}z^{r-k}$
(4.10) $=$ $\ q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{2}-q^{-2})q^{2k}z^{k}\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{x}^{+}_{\leq k-1}(q^{-2}z^{-1})$
(4.11) $=$ $\ q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}.$
@@
(4.12) $[x^{+}_{k},\boldsymbol{\phi}^{-}(z^{-1})]=$ $\ (q^{2}-1)\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{k}+(q^{4}-1)z\boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}x^{+}_{k-1}$
(4.13) $=$ $\ \boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(x^{+}_{k}+x^{+}_{k-1}z)$
(4.14) $=$ $\ \boldsymbol{\phi}^{-}(z^{-1})(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(q^{2}-1)(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{+}_{k},$
@@
$(z-q^{-2}w)\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{x}^{-}(w^{-1})=(q^{-2}z-w)\boldsymbol{x}^{-}(w^{-1})\boldsymbol{\phi}^{-}(z^{-1}).$
@@ 4.2. Basic implications of the Lu–Wang relations
Recall the notation for homogeneous components in the Drinfeld gradation introduced just after (2.14). Our next task is to compute $A_{r}^{(-1)}$, i.e., the component of degree $-1$.
Relation (2.18) (with $m=1$) implies that
@@ 4.3. $\boldsymbol{A}^{(1)}_{+}(z)$ calculation
(4.30) $\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}^{-}(z^{-1})\Omega^{-1}\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
@@
(4.36) $x^{+}_{0}x^{+}_{-1}x^{-}_{-r}=$ $\ x^{+}_{0}(x^{-}_{-r}x^{+}_{-1}-(q-q^{-1})^{-1}\phi^{-}_{-r-1})$
(4.37) $=$ $\ -(q-q^{-1})^{-1}((\phi^{-}_{-r}-\delta_{r=0}\phi^{+}_{0})x^{+}_{-1}+x^{+}_{0}\phi^{-}_{-r-1})+x^{-}_{-r}x^{+}_{0}x^{+}_{-1},$
(4.38) $x^{+}_{-1}x^{+}_{0}x^{-}_{-r}=$ $\ x^{+}_{-1}(x^{-}_{-r}x^{+}_{0}-(q-q^{-1})^{-1}(\phi^{-}_{-r}-\delta_{r=0}\phi^{+}_{0}))$
(4.39) $=$ $\ -(q-q^{-1})^{-1}(\phi^{-}_{-r-1}x^{+}_{0}+x^{+}_{-1}(\phi^{-}_{-r}-\delta_{r=0}\phi^{+}_{0}))+x^{-}_{-r}x^{+}_{-1}x^{+}_{0}$
(4.40) $=$ $\ -q^{-2}(q-q^{-1})^{-1}((\phi^{-}_{-r}-\delta_{r=0}\phi^{+}_{0})x^{+}_{-1}+x^{+}_{0}\phi^{-}_{-r-1})+x^{-}_{-r}x^{+}_{-1}x^{+}_{0},$
@@
(4.43) $\operatorname{ad}_{\overline{H}_{1}^{(2)}}\boldsymbol{A}^{(-1)}_{+}(z)=$ $\ c_{1}Kx^{+}_{-1}-q^{-2}c_{1}\sum_{r\geq 0}(q^{4}\phi^{-}_{-r-1}x^{+}_{0}+q^{6}\phi^{-}_{-r}x^{+}_{-1})z^{r}$
(4.44) $\ -c_{1}(q^{2}-q^{-2})\sum_{r\geq 0}\sum_{s=2}^{r+1}q^{2s}\phi^{-}_{s-r-1}x^{+}_{-s}z^{r}.$
(4.45) $=$ $\ c_{1}Kx^{+}_{-1}+c_{1}q^{2}K^{-1}x^{+}_{0}z^{-1}-c_{1}q^{2}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{0}z^{-1}$
(4.46) $\ -c_{1}q^{4}\boldsymbol{\phi}^{-}(z^{-1})x^{+}_{-1}-c_{1}(q^{2}-q^{-2})\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})z^{-1}.$
@@
(4.47) $\Omega\boldsymbol{A}^{(1)}_{+}(z)=$ $\ -c_{1}\boldsymbol{\phi}^{-}(z^{-1})\big{(}(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})+q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z\big{)}.$
Dividing both sides by $\Omega$ and using the fact that $\Omega$ commutes with $\boldsymbol{\phi}^{-}(z^{-1})$ completes the proof. ∎
@@
so that $\boldsymbol{A}^{(1)}_{+}(z)=\boldsymbol{\phi}^{-}(z^{-1})\mathbb{A}$.